OpenStudy (christos):

Solve 35x^2 - 125x^3 + 165x^4 = 0

OpenStudy (christos):

Please heelp?

hartnn (hartnn):

what can you factor out from all the 3 terms ?

OpenStudy (christos):

thanks for the hint <33

hartnn (hartnn):

so could you solve it ? or still need further help?

OpenStudy (christos):

I got stuck :(

hartnn (hartnn):

at which step ?

hartnn (hartnn):

35x^2 - 125x^3 + 165x^4 = 0 5x^2 (7-25x+33x^2)=0 can u proceed ? (7-25x+33x^2) =0 is a quadratic which i think you know how to solve ?

OpenStudy (anonymous):

\[35 x^2-125 x^3+165 x^4 = 0\] \[5 x^2 ( 7 -25 x + 33 x^2) = 0\]